OpenStudy (anonymous):

find the exponential function f(x)=a^x for a graph with the point (-3,8)? would i start with 8=a^-3

zepdrix (zepdrix):

Mmmmmmmmmmm, yes start there. Hmm let's see.

zepdrix (zepdrix):

\[\Large 8=\frac{1}{a^3}\]Understand what I did there?

zepdrix (zepdrix):

Property of exponents:\[\Large x^{-n}\quad=\quad \frac{1}{x^n}\]

OpenStudy (anonymous):

yeah i got that far but idk what to do next.

zepdrix (zepdrix):

Solve for a! :) Multiply both sides by a^3,\[\Large 8a^3=1\]Divide by 8,\[\Large a^3=\frac{1}{8}\]Take cube root,\[\Large a=?\]

OpenStudy (anonymous):

1/2?

zepdrix (zepdrix):

yesssss

OpenStudy (anonymous):

cool :) thanks